solve each equation The Distributive Property Fraction:
2/9 (x+18)=36

Answer by BurningFlame_marc(77)

Hello!
Distributing fraction is the same as distributing numbers. You will multiply the fraction to the numbers inside the parentheses.

Solution:
2/9 (x+18)=36
2/9(x)+4=36
2/9(x)=32
2x=288
x=144

Final Answer: 144.

I hoped you learned something from me.